Scores of an achievement test show that if follows a normal distribution. Its mean is 78 with a standard deviation of 8. Find the interval wherein the middle 80% of the scores lie?

1
Expert's answer
2021-12-07T11:37:49-0500
"\dfrac{1-0.8}{2}=0.1"

"P(X<x_1)=P(Z<\dfrac{x_1-\mu}{\sigma})"

"=P(Z<\dfrac{x_1-78}{8})=0.1"

"\dfrac{x_1-78}{8}\approx-1.281552"

"x_1=78-8(1.281552)"

"x_1\approx67.7476"

"P(X>x_2)=1-P(Z<\dfrac{x_2-\mu}{\sigma})"

"=1-P(Z<\dfrac{x_2-78}{8})=0.1"

"P(Z<\dfrac{x_2-78}{8})=0.9"

"\dfrac{x_2-78}{8}\approx1.281552"

"x_2=78+8(1.281552)"

"x_2\approx88.2524"

Interval "(67.7476, 88.2524)."